# Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit

### v2

@@ Acknowledgements.
M.M. declares that he is a Co-Founder and Shareholder of the quantum-computer company IQM Finland and of the quantum-algorithm company QMill Oy. Other authors declare no competing interests.
Viitteet
  • Kjaergaard et al. [2020] M. Kjaergaard, M. E. Schwartz, J. Braumüller, P. Krantz, J. I.-J. Wang, S. Gustavsson, and W. D. Oliver, Superconducting Qubits: Current State of Play, Annual Review of Condensed Matter Physics 11, 369 (2020).
  • Krantz et al. [2019] P. Krantz, M. Kjaergaard, F. Yan, T. P. Orlando, S. Gustavsson, and W. D. Oliver, A quantum engineer’s guide to superconducting qubits, Applied Physics Reviews 6, 021318 (2019).
  • Mamgain et al. [2023] A. Mamgain, S. S. Khaire, U. Singhal, I. Ahmad, L. A. Patel, K. D. Helambe, S. Majumder, V. Singh, and B. Suri, A Review of Developments in Superconducting Quantum Processors, Journal of the Indian Institute of Science 103, 633 (2023).
  • Huang et al. [2020] H.-L. Huang, D. Wu, D. Fan, and X. Zhu, Superconducting quantum computing: a review, Science China Information Sciences 63, 180501 (2020).
  • Ezratty [2023] O. Ezratty, Perspective on superconducting qubit quantum computing, The European Physical Journal A 59, 94 (2023).
  • Acharya et al. [2024] R. Acharya, L. Aghababaie-Beni, I. Aleiner, T. I. Andersen, M. Ansmann, F. Arute, K. Arya, A. Asfaw, N. Astrakhantsev, J. Atalaya, R. Babbush, D. Bacon, B. Ballard, J. C. Bardin, J. Bausch, A. Bengtsson, A. Bilmes, S. Blackwell, S. Boixo, G. Bortoli, A. Bourassa, J. Bovaird, L. Brill, M. Broughton, D. A. Browne, B. Buchea, B. B. Buckley, D. A. Buell, T. Burger, B. Burkett, N. Bushnell, A. Cabrera, J. Campero, H.-S. Chang, Y. Chen, Z. Chen, B. Chiaro, D. Chik, C. Chou, J. Claes, A. Y. Cleland, J. Cogan, R. Collins, P. Conner, W. Courtney, A. L. Crook, B. Curtin, S. Das, A. Davies, L. De Lorenzo, D. M. Debroy, S. Demura, M. Devoret, A. Di Paolo, P. Donohoe, I. Drozdov, A. Dunsworth, C. Earle, T. Edlich, A. Eickbusch, A. M. Elbag, M. Elzouka, C. Erickson, L. Faoro, E. Farhi, V. S. Ferreira, L. F. Burgos, E. Forati, A. G. Fowler, B. Foxen, S. Ganjam, G. Garcia, R. Gasca, E. Genois, W. Giang, C. Gidney, D. Gilboa, R. Gosula, A. G. Dau, D. Graumann, A. Greene, J. A. Gross, S. Habegger, J. Hall, M. C. Hamilton, M. Hansen, M. P. Harrigan, S. D. Harrington, F. J. H. Heras, S. Heslin, P. Heu, O. Higgott, G. Hill, J. Hilton, G. Holland, S. Hong, H.-Y. Huang, A. Huff, W. J. Huggins, L. B. Ioffe, S. V. Isakov, J. Iveland, E. Jeffrey, Z. Jiang, C. Jones, S. Jordan, C. Joshi, P. Juhas, D. Kafri, H. Kang, A. H. Karamlou, K. Kechedzhi, J. Kelly, T. Khaire, T. Khattar, M. Khezri, S. Kim, P. V. Klimov, A. R. Klots, B. Kobrin, P. Kohli, A. N. Korotkov, F. Kostritsa, R. Kothari, B. Kozlovskii, J. M. Kreikebaum, V. D. Kurilovich, N. Lacroix, D. Landhuis, T. Lange-Dei, B. W. Langley, P. Laptev, K.-M. Lau, L. L. Guevel, J. Ledford, K. Lee, Y. D. Lensky, S. Leon, B. J. Lester, W. Y. Li, Y. Li, A. T. Lill, W. Liu, W. P. Livingston, A. Locharla, E. Lucero, D. Lundahl, A. Lunt, S. Madhuk, F. D. Malone, A. Maloney, S. Mandrá, L. S. Martin, S. Martin, O. Martin, C. Maxfield, J. R. McClean, M. McEwen, S. Meeks, A. Megrant, X. Mi, K. C. Miao, A. Mieszala, R. Molavi, S. Molina, S. Montazeri, A. Morvan, R. Movassagh, W. Mruczkiewicz, O. Naaman, M. Neeley, C. Neill, A. Nersisyan, H. Neven, M. Newman, J. H. Ng, A. Nguyen, M. Nguyen, C.-H. Ni, T. E. O’Brien, W. D. Oliver, A. Opremcak, K. Ottosson, A. Petukhov, A. Pizzuto, J. Platt, R. Potter, O. Pritchard, L. P. Pryadko, C. Quintana, G. Ramachandran, M. J. Reagor, D. M. Rhodes, G. Roberts, E. Rosenberg, E. Rosenfeld, P. Roushan, N. C. Rubin, N. Saei, D. Sank, K. Sankaragomathi, K. J. Satzinger, H. F. Schurkus, C. Schuster, A. W. Senior, M. J. Shearn, A. Shorter, N. Shutty, V. Shvarts, S. Singh, V. Sivak, J. Skruzny, S. Small, V. Smelyanskiy, W. C. Smith, R. D. Somma, S. Springer, G. Sterling, D. Strain, J. Suchard, A. Szasz, A. Sztein, D. Thor, A. Torres, M. M. Torunbalci, A. Vaishnav, J. Vargas, S. Vdovichev, G. Vidal, B. Villalonga, C. V. Heidweiller, S. Waltman, S. X. Wang, B. Ware, K. Weber, T. White, K. Wong, B. W. K. Woo, C. Xing, Z. J. Yao, P. Yeh, B. Ying, J. Yoo, N. Yosri, G. Young, A. Zalcman, Y. Zhang, N. Zhu, and N. Zobrist, Quantum error correction below the surface code threshold (2024), arXiv:2408.13687 [quant-ph].
  • Koch et al. [2007] J. Koch, T. M. Yu, J. Gambetta, A. A. Houck, D. I. Schuster, J. Majer, A. Blais, M. H. Devoret, S. M. Girvin, and R. J. Schoelkopf, Charge-insensitive qubit design derived from the Cooper pair box, Physical Review A 76, 042319 (2007).
  • Blais et al. [2021] A. Blais, A. L. Grimsmo, S. Girvin, and A. Wallraff, Circuit quantum electrodynamics, Reviews of Modern Physics 93, 025005 (2021).
  • Place et al. [2021] A. P. M. Place, L. V. H. Rodgers, P. Mundada, B. M. Smitham, M. Fitzpatrick, Z. Leng, A. Premkumar, J. Bryon, A. Vrajitoarea, S. Sussman, G. Cheng, T. Madhavan, H. K. Babla, X. H. Le, Y. Gang, B. Jäck, A. Gyenis, N. Yao, R. J. Cava, N. P. De Leon, and A. A. Houck, New material platform for superconducting transmon qubits with coherence times exceeding 0.3 milliseconds, Nature Communications 12, 1779 (2021).
  • Wang et al. [2022] C. Wang, X. Li, H. Xu, Z. Li, J. Wang, Z. Yang, Z. Mi, X. Liang, T. Su, C. Yang, G. Wang, W. Wang, Y. Li, M. Chen, C. Li, K. Linghu, J. Han, Y. Zhang, Y. Feng, Y. Song, T. Ma, J. Zhang, R. Wang, P. Zhao, W. Liu, G. Xue, Y. Jin, and H. Yu, Towards practical quantum computers: transmon qubit with a lifetime approaching 0.5 milliseconds, npj Quantum Information 8, 3 (2022).
  • Gordon et al. [2022] R. T. Gordon, C. E. Murray, C. Kurter, M. Sandberg, S. A. Hall, K. Balakrishnan, R. Shelby, B. Wacaser, A. A. Stabile, J. W. Sleight, M. Brink, M. B. Rothwell, K. P. Rodbell, O. Dial, and M. Steffen, Environmental radiation impact on lifetimes and quasiparticle tunneling rates of fixed-frequency transmon qubits, Applied Physics Letters 120, 074002 (2022).
  • Deng et al. [2023] H. Deng, Z. Song, R. Gao, T. Xia, F. Bao, X. Jiang, H.-S. Ku, Z. Li, X. Ma, J. Qin, H. Sun, C. Tang, T. Wang, F. Wu, W. Yu, G. Zhang, X. Zhang, J. Zhou, X. Zhu, Y. Shi, H.-H. Zhao, and C. Deng, Titanium Nitride Film on Sapphire Substrate with Low Dielectric Loss for Superconducting Qubits, Physical Review Applied 19, 024013 (2023).
  • Sivak et al. [2023] V. V. Sivak, A. Eickbusch, B. Royer, S. Singh, I. Tsioutsios, S. Ganjam, A. Miano, B. L. Brock, A. Z. Ding, L. Frunzio, S. M. Girvin, R. J. Schoelkopf, and M. H. Devoret, Real-time quantum error correction beyond break-even, Nature 616, 50 (2023).
  • Biznárová et al. [2023] J. Biznárová, A. Osman, E. Rehnman, L. Chayanun, C. Križan, P. Malmberg, M. Rommel, C. Warren, P. Delsing, A. Yurgens, J. Bylander, and A. F. Roudsari, Mitigation of interfacial dielectric loss in aluminum-on-silicon superconducting qubits (2023), arXiv:2310.06797 [cond-mat, physics:quant-ph].
  • Bal et al. [2024] M. Bal, A. A. Murthy, S. Zhu, F. Crisa, X. You, Z. Huang, T. Roy, J. Lee, D. V. Zanten, R. Pilipenko, I. Nekrashevich, A. Lunin, D. Bafia, Y. Krasnikova, C. J. Kopas, E. O. Lachman, D. Miller, J. Y. Mutus, M. J. Reagor, H. Cansizoglu, J. Marshall, D. P. Pappas, K. Vu, K. Yadavalli, J.-S. Oh, L. Zhou, M. J. Kramer, F. Lecocq, D. P. Goronzy, C. G. Torres-Castanedo, P. G. Pritchard, V. P. Dravid, J. M. Rondinelli, M. J. Bedzyk, M. C. Hersam, J. Zasadzinski, J. Koch, J. A. Sauls, A. Romanenko, and A. Grassellino, Systematic improvements in transmon qubit coherence enabled by niobium surface encapsulation, npj Quantum Information 10, 43 (2024).
  • Kono et al. [2024] S. Kono, J. Pan, M. Chegnizadeh, X. Wang, A. Youssefi, M. Scigliuzzo, and T. J. Kippenberg, Mechanically induced correlated errors on superconducting qubits with relaxation times exceeding 0.4 ms, Nature Communications 15, 3950 (2024).
  • Dunsworth et al. [2017] A. Dunsworth, A. Megrant, C. Quintana, Z. Chen, R. Barends, B. Burkett, B. Foxen, Y. Chen, B. Chiaro, A. Fowler, R. Graff, E. Jeffrey, J. Kelly, E. Lucero, J. Y. Mutus, M. Neeley, C. Neill, P. Roushan, D. Sank, A. Vainsencher, J. Wenner, T. C. White, and J. M. Martinis, Characterization and reduction of capacitive loss induced by sub-micron Josephson junction fabrication in superconducting qubits, Applied Physics Letters 111, 022601 (2017).
  • Wang et al. [2019] Z. Wang, S. Shankar, Z. Minev, P. Campagne-Ibarcq, A. Narla, and M. Devoret, Cavity Attenuators for Superconducting Qubits, Physical Review Applied 11, 014031 (2019).
  • Nersisyan et al. [2019] A. Nersisyan, E. A. Sete, S. Stanwyck, A. Bestwick, M. Reagor, S. Poletto, N. Alidoust, R. Manenti, R. Renzas, C.-V. Bui, K. Vu, T. Whyland, and Y. Mohan, Manufacturing low dissipation superconducting quantum processors, in 2019 IEEE International Electron Devices Meeting (IEDM) (IEEE, San Francisco, CA, USA, 2019) pp. 31.1.1–31.1.4.
  • Wei et al. [2020] K. X. Wei, I. Lauer, S. Srinivasan, N. Sundaresan, D. T. McClure, D. Toyli, D. C. McKay, J. M. Gambetta, and S. Sheldon, Verifying multipartite entangled Greenberger-Horne-Zeilinger states via multiple quantum coherences, Physical Review A 101, 032343 (2020).
  • Moskalev et al. [2023] D. O. Moskalev, E. V. Zikiy, A. A. Pishchimova, D. A. Ezenkova, N. S. Smirnov, A. I. Ivanov, N. D. Korshakov, and I. A. Rodionov, Optimization of shadow evaporation and oxidation for reproducible quantum Josephson junction circuits, Scientific Reports 13, 4174 (2023).
  • Pishchimova et al. [2023] A. A. Pishchimova, N. S. Smirnov, D. A. Ezenkova, E. A. Krivko, E. V. Zikiy, D. O. Moskalev, A. I. Ivanov, N. D. Korshakov, and I. A. Rodionov, Improving Josephson junction reproducibility for superconducting quantum circuits: junction area fluctuation, Scientific Reports 13, 6772 (2023).
  • Pop et al. [2014] I. M. Pop, K. Geerlings, G. Catelani, R. J. Schoelkopf, L. I. Glazman, and M. H. Devoret, Coherent suppression of electromagnetic dissipation due to superconducting quasiparticles, Nature 508, 369 (2014).
  • Gyenis et al. [2021] A. Gyenis, P. S. Mundada, A. Di Paolo, T. M. Hazard, X. You, D. I. Schuster, J. Koch, A. Blais, and A. A. Houck, Experimental realization of an intrinsically error-protected superconducting qubit, PRX Quantum 2, 010339 (2021), arXiv:1910.07542 [cond-mat, physics:quant-ph].
  • Somoroff et al. [2023] A. Somoroff, Q. Ficheux, R. A. Mencia, H. Xiong, R. Kuzmin, and V. E. Manucharyan, Millisecond Coherence in a Superconducting Qubit, Physical Review Letters 130, 267001 (2023).
  • Jeffrey et al. [2014] E. Jeffrey, D. Sank, J. Mutus, T. White, J. Kelly, R. Barends, Y. Chen, Z. Chen, B. Chiaro, A. Dunsworth, A. Megrant, P. O’Malley, C. Neill, P. Roushan, A. Vainsencher, J. Wenner, A. Cleland, and J. M. Martinis, Fast Accurate State Measurement with Superconducting Qubits, Physical Review Letters 112, 190504 (2014).
  • Sete et al. [2015] E. A. Sete, J. M. Martinis, and A. N. Korotkov, Quantum theory of a bandpass Purcell filter for qubit readout, Physical Review A 92, 012325 (2015).
  • Tuokkola et al. [2024] M. Tuokkola, Y. Sunada, H. Kivijärvi, J. Albanese, L. Grönberg, J.-P. Kaikkonen, V. Vesterinen, J. Govenius, and M. Möttönen, Data and GDS file for ”Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit”missing (2024).
  • Burnett et al. [2018] J. Burnett, A. Bengtsson, D. Niepce, and J. Bylander, Noise and loss of superconducting aluminium resonators at single photon energies, Journal of Physics: Conference Series 969, 012131 (2018).
  • Verjauw et al. [2021] J. Verjauw, A. Potočnik, M. Mongillo, R. Acharya, F. Mohiyaddin, G. Simion, A. Pacco, T. Ivanov, D. Wan, A. Vanleenhove, L. Souriau, J. Jussot, A. Thiam, J. Swerts, X. Piao, S. Couet, M. Heyns, B. Govoreanu, and I. Radu, Investigation of Microwave Loss Induced by Oxide Regrowth in High- Q Niobium Resonators, Physical Review Applied 16, 014018 (2021).
  • Osman et al. [2021] A. Osman, J. Simon, A. Bengtsson, S. Kosen, P. Krantz, D. P. Lozano, M. Scigliuzzo, P. Delsing, J. Bylander, and A. Fadavi Roudsari, Simplified Josephson-junction fabrication process for reproducibly high-performance superconducting qubits, Applied Physics Letters 118, 064002 (2021).
  • Bilmes et al. [2021] A. Bilmes, A. K. Händel, S. Volosheniuk, A. V. Ustinov, and J. Lisenfeld, In-situ bandaged Josephson junctions for superconducting quantum processors, Superconductor Science and Technology 34, 125011 (2021).
  • Stefanazzi et al. [2022] L. Stefanazzi, K. Treptow, N. Wilcer, C. Stoughton, C. Bradford, S. Uemura, S. Zorzetti, S. Montella, G. Cancelo, S. Sussman, A. Houck, S. Saxena, H. Arnaldi, A. Agrawal, H. Zhang, C. Ding, and D. I. Schuster, The QICK (Quantum Instrumentation Control Kit): Readout and control for qubits and detectors, Review of Scientific Instruments 93, 044709 (2022).
  • Ding et al. [2024] C. Ding, M. Di Federico, M. Hatridge, A. Houck, S. Leger, J. Martinez, C. Miao, D. S. I, L. Stefanazzi, C. Stoughton, S. Sussman, K. Treptow, S. Uemura, N. Wilcer, H. Zhang, C. Zhou, and G. Cancelo, Experimental advances with the QICK (Quantum Instrumentation Control Kit) for superconducting quantum hardware, Physical Review Research 6, 013305 (2024).
  • Hyyppä et al. [2022] E. Hyyppä, S. Kundu, C. F. Chan, A. Gunyhó, J. Hotari, D. Janzso, K. Juliusson, O. Kiuru, J. Kotilahti, A. Landra, W. Liu, F. Marxer, A. Mäkinen, J.-L. Orgiazzi, M. Palma, M. Savytskyi, F. Tosto, J. Tuorila, V. Vadimov, T. Li, C. Ockeloen-Korppi, J. Heinsoo, K. Y. Tan, J. Hassel, and M. Möttönen, Unimon qubit, Nature Communications 13, 6895 (2022).
Liite A Extended Data Figures and Tables
@@
9 3.73 303 7.1 201
10 3.92 401 9.9 N/A
11 3.86 393 9.5 N/A
12 4.24 292 7.7 N/A
13 5.92 280 10.4 238
14 3.02 270 5.1 307
15 3.95 323 8.0 286
16 4.79 210 6.5 290
This work 2.89 502 9.1 541
@@ 80 Plus
Low pass filters RLC F-30-8000-R (3-dB point = 8.4 GHz, 60-dB stopband = 10.8–40 GHz) LPF
@@
Extended Data Fig. 3: Energy relaxation and echo dephasing times and their relative uncertainties for qubit $\bm{\mathrm{Q}_{2}}$ measured during the second cooldown. (a) $T_{1}$ measured while pumping the TWPA. (b) $T_{1}$ measured without pumping the TWPA. (c) $T_{2}^{\textrm{echo}}$ measured while pumping the TWPA. (d) $T_{2}^{\textrm{echo}}$ measured without pumping the TWPA. The orange vertical lines and the pink horizontal lines represent the average uncertainties and the median values of $T_{1}$ or $T_{2}^{\textrm{echo}}$, respectively. The values in the grey area are excluded from further analysis.

### v3

@@ Acknowledgements.
M.M. declares that he is a Co-Founder and Shareholder of the quantum-computer company IQM Finland and of the quantum-algorithm company QMill Oy. J.G and V.V declare that they are Co-Founders and Shareholders of the quantum-hardware company Arctic Instruments Oy. Other authors declare no competing interests.
Appendix A Extended Data Figures and Tables
@@
place_new_2021 3.73 303 7.1 201
wang_towards_2022 3.92 401 9.9 N/A
gordon_environmental_2022 3.86 393 9.5 N/A
deng_titanium_2023 4.24 292 7.7 N/A
sivak_real-time_2023 5.92 280 10.4 238
biznarova_mitigation_2023 3.02 270 5.1 307
bal_systematic_2024 3.95 323 8.0 286
kono_mechanically_2024 4.79 210 6.5 290
This work 2.86 425 7.7222Qubit frequency in the relaxation time measurement is 2.89 GHz. 541
@@ 80 Plus
Low-pass filters RLC F-30-8000-R (3-dB point = 8.4 GHz, 60-dB stopband = 10.8–40 GHz) LPF
@@
Extended Data Fig. 3: Energy relaxation and echo dephasing times and their relative uncertainties for qubit $\bm{\mathrm{Q}_{2}}$ measured during the second cooldown. (a) $T_{1}$ measured while pumping the TWPA. (b) $T_{1}$ measured without pumping the TWPA. (c) $T_{2}^{\textrm{echo}}$ measured while pumping the TWPA. (d) $T_{2}^{\textrm{echo}}$ measured without pumping the TWPA. The orange horizontal lines and the pink vertical lines represent the average uncertainties and the median values of $T_{1}$ or $T_{2}^{\textrm{echo}}$, respectively. The values in the grey area are excluded from further analysis.
@@
Supplementary Information for ”Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit”
Supplementary Note S1: High-coherence transmon qubit in a 3D cavity
In this supplementary section, we present high-coherence transmon qubits coupled to three-dimensional (3D) cavity resonators to demonstrate the reproducibility of the fabrication process described in the main article. We measure two additional qubits $\mathrm{Q}_{\text{S}1}$ and $\mathrm{Q}_{\text{S}2}$ fabricated with an almost identical recipe as described in Sec. II. The only differences are the HF etching time which is 5 min instead of 10 min, and the junction oxidation pressure which is 1.5 mbar instead of 1.2 mbar. The qubits have an identical design, which is shown in Supplementary Fig. S1 and is available on Zenodo [1].
Qubits $\mathrm{Q}_{\text{S}1}$ and $\mathrm{Q}_{\text{S}2}$ are measured in 3D cavity resonators, C20 and C16, respectively. Supplementary Fig. S2 presents illustrations of the cavities, which are made of 1050 aluminum alloy, and the design is available on Zenodo [1].
(a)(b)(c)
Supplementary Fig. S1: Supplementary sample design. (a) Design of the transmon samples used in the measurements of this Supplementary Note S1 and (b, c) magnifications in the vicinity of the Josephson junction. See full design file in Zenodo.
(a)(b)
Supplementary Fig. S2: Supplementary cavity design. (a, b) Design illustrations of the 3D aluminum cavities (a) C16 and (b) C20 used in the experiments of this supplementary note. See full design file in Zenodo.
The qubits are measured in the same dilution refrigerator and with the same evaluation board as in the main article. Unlike in the main article, in this measurement setup, the readout signal and the qubit drive signal are combined into a single line by using a directional coupler, and we measure the reflection from the cavity instead of the transmission. The measurement setup is presented in the schematic diagram of Supplementary Fig. S3, and the equipment is listed in Tables 3 and S1. The cavity is protected from electromagnetic fields with a custom-designed four-layer magnetic shield. From the innermost to the outermost layer, the shield consists of 1.5-mm-thick gold-plated copper (C10100), 1.5-mm-thick aluminum 1050, and two layers of 1.5-mm-thick mumetal (Hymu80).
We measure that the cavities C20 and C16 have resonance frequencies of 5.55 GHz and 6.32 GHz, respectively. The frequency of qubit $\mathrm{Q}_{\text{S}1}$ is measured to be 2.9755 GHz and with an anharmonicity of $-226$ MHz. These parameters are 2.694 GHz and $-230$ MHz for the qubit $\mathrm{Q}_{\text{S}2}$, respectively. We measure the qubit $\mathrm{Q}_{\text{S}1}$ to have a median energy relaxation time $T_{1}$ of $270$ µs with the longest measured $T_{1}$ of $(521\pm 27)$ µs. For the qubit $\mathrm{Q}_{\text{S}2}$, the median $T_{1}$ is measured to be $288$ µs and the longest $T_{1}$ to be $(496\pm 26)$ µs. Supplementary Fig. S4 presents distributions of the measured relaxation times and the time traces of the longest measured relaxation times. The data used to generate the figures are published on Zenodo [1].
Supplementary Fig. S3: Schematic of the measurement setup used in measurements of the qubits in cavities. See Extended Data Table 3 and Supplementary Table S1 for a description of each component.
Supplementary Table S1: Equipment used in the measurement setup of the supplementary note.
Equipment Description Abbreviation
Low-pass filters Mini-Circuits VLF-3400+ (3-dB point = 3.8 GHz) LPF1
Low-noise amplifier Mini-Circuits ZX60-123LN-S+ LN
Directional coupler Mini-Circuits ZUDC20-02183-S+
High-pass filter Mini-Circuits VHF-3100+ (20-dB stopband = DC–2.5 GHz) HPF
Mini-Circuits VHF-4400+ (20-dB stopband = DC–3.85 GHz) HPF
Eccosorb filter Kawashima Manufacturing 3.5mm Filter Connector L5 (0.32 dB/GHz) Ecco.
Circulator Low Noise Factory LNF-CIC4_8A (4–8 GHz)
(a)(b)(c)(d)
Supplementary Fig. S4: Energy relaxation times of the additional qubits $\bm{\mathrm{Q}_{\text{S}1}}$ and $\bm{\mathrm{Q}_{\text{S}2}}$. (a, b) Distributions of the measured energy relaxation times of the qubits (a)  $\mathrm{Q}_{\text{S}1}$ over 48 hours and (b)  $\mathrm{Q}_{\text{S}2}$ over 40 hours. (c, d) Time traces for the longest measured relaxation times of (c) the qubit $\mathrm{Q}_{\text{S}1}$ with $\mathrm{T}_{1}=(521\pm 27)$ µs and (d) the qubit $\mathrm{Q}_{\text{S}2}$ with $\mathrm{T}_{1}=(496\pm 26)$ µs.